# Tactile-based Object Retrieval From Granular Media

### v2

@@ Abstract
We introduce GEOTACT, a robotic manipulation method capable of retrieving objects buried in granular media. This is a challenging task due to the need to interact with granular media, and doing so based exclusively on tactile feedback, since a buried object can be completely hidden from vision. Tactile feedback is in itself challenging in this context, due to ubiquitous contact with the surrounding media, and the inherent noise level induced by the tactile readings. To address these challenges, we use a learning method trained end-to-end with simulated sensor noise. We show that our problem formulation leads to the natural emergence of learned pushing behaviors that the manipulator uses to reduce uncertainty and funnel the object to a stable grasp despite spurious and noisy tactile readings. We also introduce a training curriculum that enables learning these behaviors in simulation, followed by zero-shot transfer to real hardware. To the best of our knowledge, GEOTACT is the first method to reliably retrieve a number of different objects from a granular environment, doing so on real hardware and with integrated tactile sensing.
Index Terms:
@@
I Introduction
Interaction with or retrieval of objects buried in granular media (such as sand, pebbles, etc.) is important to a wide range of applications, including search and rescue, archaeology, demining, seabed exploration, etc. Humans are experts at performing such tasks, using only tactile sensing to compensate for the absence of visual feedback. However, transferring such behaviors to robots poses multiple challenges.
@@ I Introduction
Second, manipulation within granular media places specific requirements on the hardware: streamlined shapes are important for smooth movement, and contacts come from every angle and direction, thus requiring extensive sensing coverage. Despite many breakthroughs [1, 2, 3, 4, 5, 6] in sensor developments, many sensors are either flat or do not have sensing coverage around edges and corners. Due to such compounding limitations, robotic manipulation in granular media remains poorly explored compared to tactile-based work with objects in open-air environments.
@@
In this work, we propose GEOTACT (Granular Environment Object Tracing and ACquisition via Touch), a novel method for retrieving objects buried under granular media. GEOTACT relies exclusively on tactile sensing, as we assume the object can be completely hidden from any vision sensors. In particular, we use the DISCO tactile finger [7], which provides both the streamlined shape and sensing coverage needed to enable the task.
@@
Critically, GEOTACT relies on a pushing behavior which further reduces uncertainty and increases robustness against noisy tactile readings. Despite pushing having been shown in the literature to be a key element of many manipulator operations [8, 9, 10, 11], no previous work has utilized pushing action to grasp objects within granular media. The GEOTACT formulation of the learning problem action space leads to the emergence of a series of pushing behaviors that funnels the target object into a stable grasp.
To enable this learning process, we propose a curriculum strategy that enables learning in simulation, followed by zero-shot transfer to real hardware. We first pre-train our policy in an open-air environment and then fine-tune it inside simulated granular media. Despite the recent effort from the community to improve granular material simulation accuracy and efficiency [12, 13, 14, 15, 16, 17, 18, 19, 20], granular media simulation are still significantly slower than open-air simulation. However, our curriculum strategy helps to bootstrap this training process, allowing us to reach a better policy within a shorter training time by transferring the skills learned from the open air to granular media.
@@
  • To the best of our knowledge, we are the first to demonstrate robotic grasping of objects buried under granular media using only touch sensing.
@@
Our quantitative real-robot experiments show that GEOTACT consistently achieves success rates higher than 60% on a majority of objects from both its training set (seven objects), and testing set (six additional objects not seen during training). In additional testing, we also show promising early results on an extended set of 22 objects, including rigid, deformable, and articulated objects with various complex shapes. Our extensive experiments in both simulation and on the real robot show that our method is more robust to sensor noise and achieves a higher success rate compared to baselines.
II Related Work
II-A Robot Interaction with Granular Media
@@
Unlike manipulating objects buried inside granular media, there are many more works focusing on manipulating granular materials directly, such as scooping and bulldozing [21, 13, 22], plate dragging (soil-tool interaction) [23, 24, 25], and pouring [26, 27, 28, 29]. In addition to tactile feedback, some works [30, 31] also incorporate visual and audio feedback for scooping and pouring tasks. Other works [32, 33, 34] study legged locomotion on granular materials. In terms of hardware design, there are works that use granular materials and particle jamming to design gripper [35, 36], manipulator [37], swarm robots [38], or haptic displays [39, 40].
@@
In contrast, interacting with objects buried inside granular materials has only recently begun receiving attention from the research community, but is only limited to tactile perception such as object mapping, localization or identification. Jia, et al. [41] teleoperate a robotic hand with the BioTac tactile sensors to detect contact with a cylinder fixed inside a bed of granular media. They take advantage of the multi-modal sensors to distinguish between contact with an object embedded in granular media and ubiquitous contact with surrounding granular media. In their follow-up work [42], instead of teleoperation, they develop a tactile exploration policy for efficient localization and mapping of the buried object. Patel, et al. [1] develop a new tactile finger built on top of GelSight technology [3, 43] that facilitates penetration in granular media with the help of mechanical vibrations. They use this sensor to identify four simple 2D shapes i.e. triangle, square, hexagon, and circle. More recently, Zhang, et al. [44] propose a proximity sensing system to detect an object’s presence without contact and finally use a Bayesian-optimization-algorithm-guided exploration strategy to localize buried objects. Different from all these aforementioned works, GEOTACT goes beyond tactile perception and focuses on retrieving any buried objects. To the best of our knowledge, ours is the first work to enable a robotic gripper to retrieve buried objects from granular media.
II-B Robotic Pushing for Grasping
@@
Pushing is a key element of many manipulator operations. It is a good way to reduce uncertainty in the locations of objects, to move many objects at once, and to move objects that are hard to grasp. Combining both non-prehensile (e.g., pushing) and prehensile (e.g., grasping) manipulation policies has received increasing attention from the community. Brost [8] presents an algorithm that plans parallel-jaw push-grasping motions for polygonal objects with pose uncertainty, where the object is pushed by one plate towards the second one, and then squeezed between the two. Mason [9] investigates the mechanics and planning of pushing in object manipulation under uncertainty. The seminal work of Dogar et al. [10, 11] presents a robust planning framework for push-grasping (non-prehensile motions baked within grasping primitives) to reduce grasp uncertainty.
@@
While the policies in the works above remain largely hand-crafted, other methods [45, 46] explore the model-free planning of pushing motions to move objects to target positions that are more favorable for pre-designed grasping algorithms. Boularias et al. [47] explores the use of reinforcement learning for training control policies to select among push and grasp proposals represented by hand-crafted features. However, their method models perception and control policies separately (not end-to-end); it relies on model-based simulation to predict the motion of pushed objects and to infer its benefits for future grasping. Similarly, Yu et al. [48] also separates push planning and grasp planning by training two networks. More recently, some works [49, 50, 51, 52] have been training model-free end-to-end policy using reinforcement learning to push and separate objects in clutter for better grasping. Our method is also model-free and trained end-to-end, making no assumptions about the shapes or dynamics of the objects. However, unlike all the previous works cited here, ours is the first to demonstrate a pushing behavior for grasping inside granular media.
II-C Active Tactile Exploration
@@
Tactile sensors need an intelligent exploration policy to actively interact with the target object and efficiently gather sparse and local tactile readings. These exploration policies can be roughly divided into heuristics-based and learning-based. One of the most popular heuristics-based exploration policies is contour-following [53, 54, 55]. Other heuristics include information gain (uncertainty reduction) [56, 57, 58, 59], attention cubes [60], Monte Carlo tree search [61], and dynamic potential fields [62]. Some works study tactile exploration policy specifically for decision making, such as object identification, where they pre-train a discriminator with a pre-collected dataset and then use it to estimate action quality with Bayesian methods to reduce uncertainty [63, 64, 65, 66, 67]. Depending on the task, heuristic-based exploration policies can improve efficiency and require no training, but their performance can deteriorate significantly when sensor noise shows up.
@@
In contrast, Xu et al. [68, 69] study learning-based exploration policy trained through reinforcement learning in an unsupervised fashion. They also aim for object identification, and their method has two separate components, trained jointly: a discriminator for decision-making and an explorer for guiding the exploration. Different from these previous works, our method aims to grasp buried objects inside granular media, which is a more challenging and novel task. In addition, our method only has one end-to-end policy that guides both the pushing and grasping actions.
III Method
@@
We formulate our problem as follows. A robotic manipulator is presented with a workspace filled with granular media. An object is buried completely under the granular media, placed such that the gripper can contact it by moving forward, but with unknown orientation or exact position. The goal of the manipulator is to retrieve the buried object, by lifting it clear of the granular environment. We assume that the object can start out fully buried with no part visible from the surface. Our goal is to perform the task without relying on vision and using only touch sensing.
@@
Tactile sensing in granular media requires the ability to distinguish between readings due to ubiquitous contacts with the granules themselves, and readings indicating contact with the target object. As explained later in this section, we use a simple filtering technique to make this distinction. However, the nature of operation in such environments increases the noise inherent in any tactile sensor. We thus need an exploration policy that is well suited for handling tactile noise and spurious or erroneous readings.
@@
III-A Action Space Definition and Pushing Behavior
We formulate our problem as a model-free reinforcement learning task. Object exploration and retrieval are achieved through a series of discrete translation/rotation actions of the gripper. The episode starts after the finger obtains the first contact. At each step, our policy takes in a sequence of contact locations and forces from the past steps and then outputs the next action for the gripper to take for exploration. Our policy also determines when to stop pushing and execute the grasp.
@@ III-A Action Space Definition and Pushing Behavior
Specifically, our action primitives consist of a planar pushing action and a grasping action. The gripper moves in a 2D plane and is always perpendicular to the table, as shown in Fig. 1. Given the current gripper pose (position $[x,y]$ and rotation $\theta$), the action is a 4-dimensional vector $[dx,dy,d\theta,T]$. A value of $T=0$ indicates a pushing action, in which case $dx,dy,d\theta$ indicate the change to $x$, $y$, and $\theta$, respectively. For $x,y$, the change is 1cm and for $\theta$, the change is $15^{\circ}$; namely, $dx,dy\in\{-1\text{cm},0,+1\text{cm}\}$ and $d\theta\in\{-15^{\circ},0,+15^{\circ}\}$. Notice that in this action space formulation, it is possible to change $x,y$, and $\theta$ simultaneously in one action. $T=1$ corresponds to a grasping action. In this case, the values of $dx,dy,d\theta$ are ignored, the gripper is closed, and the object is lifted. The episode is terminated when a grasping action is executed or a predefined maximum number of steps is reached.
As discussed in the results section, this formulation results in the learned policy effectively making use of sequences of pushing behaviors of various lengths, depending on the nature of observations received during execution. However, a critical aspect of learning such behaviors is training in the presence of noise in the policy observation space, which we discuss next.
III-B Observation Space and Tactile Data Processing
@@
The observation of our policy is a sequence of contact locations, net forces, and previous actions from the last ten steps. For each step, the observation contains one contact location and one net force for each finger, as well as the action from the previous step. Both the contact location and the net force are 3-dimensional vectors under the finger’s frame of reference (using the center of the finger base as the origin). The previous action is a 4-dimensional vector. When there is no contact recorded after a particular step, we simply fill in zeros in the location and force vectors. Compared to high-dimensional or multimodal tactile data, contact locations and forces are easy to simulate, enabling a large amount of training in simulation and zero-shot transfer to real-world experiments.
@@
Tactile observation in granular media is much more noisy compared to open air. First, there might not be direct contact between the finger and the target object when pushing contacts are detected. Instead, there could be a layer of granules between the finger and the target object. This layer can potentially introduce more noise to the force directions. Second, for small and light objects, pushing contact can be hard to distinguish from ubiquitous contacts, and, depending on how fast the finger accelerates and moves inside the medium, pushing contacts can be falsely reported even though the finger has not started pushing the target object. Ideally, object- and granular media-specific force filtering thresholds might be needed for better filtering accuracy. However, in this we use a fixed filtering value for all objects in both simulation and on the real hardware, and observe that our policy learns to be robust to spurious pushing contacts.
III-C Reward Design and Policy Architecture

### v3

@@ Abstract
We introduce GEOTACT, the first robotic system capable of grasping and retrieving objects of potentially unknown shapes buried in a granular environment. While important in many applications, ranging from mining and exploration to search and rescue, this type of interaction with granular media is difficult due to the uncertainty stemming from visual occlusion and noisy contact signals. To address these challenges, we use a learning method relying exclusively on touch feedback, trained end-to-end with simulated sensor noise. We show that our problem formulation leads to the natural emergence of learned pushing behaviors that the manipulator uses to reduce uncertainty and funnel the object to a stable grasp despite spurious and noisy tactile readings. We introduce a training curriculum that bootstraps learning in simulated granular environments, enabling zero-shot transfer to real hardware. Despite being trained only on seven objects with primitive shapes, our method is shown to successfully retrieve 35 different objects, including rigid, deformable, and articulated objects with complex shapes.
keywords:
@@
1 Introduction
Interacting with granular media, for example, in order to retrieve buried objects, is a foundational skill for numerous important robotics applications. Examples include mining and demining, search and rescue, excavation and construction, archaeology, and exploration of terrestrial, seabed, and even extraterrestrial environments, all requiring robots that can extract valuable minerals, metals, or other geological materials from the ground.
Many of these aforementioned applications require operation in hazardous environments that pose a significant danger to human workers. For example, mining or rescue sites are often located in remote and inhospitable areas. In the United States, about 200 casualties occur per year from cave-ins, ground collapses, or other excavation incidents alone [lew2002excavation]. Looking further out, future extraterrestrial sites that are currently being contemplated, such as the moon [creech2022artemis] or even Mars [zhang2022analysis, nasamars], could completely prevent on-site operators. As a result, enabling autonomous robots to replace humans in searching through loose soil or other granular media could play an important role in reducing exposure to injuries. However, while humans are experts at performing such tasks, for example, using tactile sensing to compensate when visual feedback is absent, transferring such behaviors to robots poses multiple challenges.
@@ 1 Introduction
Second, manipulation within granular media places specific requirements on the hardware. In a granular environment, contacts come from every angle and direction, thus requiring extensive sensing coverage. Despite many breakthroughs [patel2021digger, lambeta2020digit, johnson2009retrographic, yuan2017gelsight, dong2017improved, chang2023investigation] in sensor developments, many sensors are either flat or do not have sensing coverage around edges and corners. In addition, we want an intruder shape that can render low resistance and smooth movement inside granular media. For example, circular and elliptical shapes typically experience slightly less resistance during horizontal dragging and significantly less resistance during vertical lifting compared to rectangular shapes [tripura2022role]. Due to such compounding limitations, robotic manipulation in granular media remains poorly explored compared to tactile-based work with objects in open-air environments.
@@
In this work, we propose GEOTACT (Granular Environment Object Tracing and ACquisition via Touch), a novel method for retrieving objects buried under granular media. GEOTACT relies exclusively on tactile sensing, as we assume the object can be completely hidden from any vision sensors. In particular, we use the DISCO tactile finger [piacenza2020sensorized], which provides both the streamlined shape and sensing coverage needed to enable the task.
@@
Critically, GEOTACT relies on a pushing behavior which further reduces uncertainty and increases robustness against noisy tactile readings. Despite pushing having been shown in the literature to be a key element of many manipulator operations [brost1988automatic, mason1986mechanics, dogar2010push, dogar2012planning], no previous work has utilized pushing action to grasp objects within granular media. The GEOTACT formulation of the learning problem action space leads to the emergence of a series of pushing behaviors that funnels the target object into a stable grasp.
To enable this learning process, we propose a curriculum strategy that enables learning in simulation, followed by zero-shot transfer to real hardware. We first pre-train our policy in an open-air environment and then fine-tune it inside simulated granular media. Despite the recent effort from the community to improve granular material simulation accuracy and efficiency [kloss2012models, millard2023granular, thompson2022lammps, xian2023fluidlab, fang2021chrono, servin2021multiscale, ansys, haeri2020efficient, hu2019difftaichi], granular media simulation are still significantly slower than open-air simulation. However, our curriculum strategy helps to bootstrap this training process, allowing us to reach a better policy within a shorter training time by transferring the skills learned from the open air to granular media.
@@
  • To the best of our knowledge, we are the first to demonstrate robotic grasping of objects buried under a specific type of granular media using only touch sensing.
@@
Our quantitative real-robot experiments show that GEOTACT achieves success rates of 54% and 63% on its training set (seven objects) and testing set (six additional objects not seen during training), respectively. In additional testing, we also show promising early results on an extended set of 22 objects, including rigid, deformable, and articulated objects with various complex shapes. Our extensive experiments in both simulation and on the real robot show that our method is more robust to sensor noise and achieves a higher success rate compared to baselines.
2 Related Work
@@
Robot Interaction with Granular Media. Unlike manipulating objects buried inside granular media, there are many more works focusing on manipulating granular materials directly, such as scooping and bulldozing [sarata2004trajectory, millard2023granular, zhu2023few], plate dragging (soil-tool interaction) [kobayakawa2020interaction, swick1988model, gravish2014force], and pouring [wu2020squirl, yamaguchi2015pouring, matl2020inferring, tuomainen2022manipulation]. In addition to tactile feedback, some works [clarke2018learning, clarke2019robot] also incorporate visual and audio feedback for scooping and pouring tasks. Other works [li2013terradynamics, hauser2016friction, zhu2019data] study legged locomotion on granular materials. In terms of hardware design, there are works that use granular materials and particle jamming to design gripper [brown2010universal, fakhri2023systematic], manipulator [cheng2012design], swarm robots [karimi2020boundary], or haptic displays [brown2020soft, stanley2013haptic].
@@
In contrast, interacting with objects buried inside granular materials has only recently begun receiving attention from the research community, but is only limited to tactile perception such as object mapping, localization or identification. Jia, et al. [jia2021tactile] teleoperate a robotic hand with the BioTac tactile sensors to detect contact with a cylinder fixed inside a bed of granular media. They take advantage of the multi-modal sensors to distinguish between contact with an object embedded in granular media and ubiquitous contact with surrounding granular media. In their follow-up work [jia2022autonomous], instead of teleoperation, they develop a tactile exploration policy for efficient localization and mapping of the buried object. Patel, et al. [patel2021digger] develop a new tactile finger built on top of GelSight technology [johnson2009retrographic, johnson2011microgeometry] that facilitates penetration in granular media with the help of mechanical vibrations. They use this sensor to identify four simple 2D shapes i.e. triangle, square, hexagon, and circle. More recently, Zhang, et al. [zhang2023grains] propose a proximity sensing system to detect an object’s presence without contact and finally use a Bayesian-optimization-algorithm-guided exploration strategy to localize buried objects. Different from all these aforementioned works, GEOTACT goes beyond tactile perception and focuses on retrieving any buried objects. To the best of our knowledge, ours is the first work to enable a robotic gripper to retrieve buried objects from granular media.
@@
Robotic Pushing for Grasping. Pushing is a key element of many manipulator operations. It is a good way to reduce uncertainty in the locations of objects, to move many objects at once, and to move objects that are hard to grasp. Combining both non-prehensile (e.g., pushing) and prehensile (e.g., grasping) manipulation policies has received increasing attention from the community. Brost [brost1988automatic] presents an algorithm that plans parallel-jaw push-grasping motions for polygonal objects with pose uncertainty, where the object is pushed by one plate towards the second one, and then squeezed between the two. Mason [mason1986mechanics] investigates the mechanics and planning of pushing in object manipulation under uncertainty. The seminal work of Dogar et al. [dogar2010push, dogar2012planning] presents a robust planning framework for push-grasping (non-prehensile motions baked within grasping primitives) to reduce grasp uncertainty.
@@
While the policies in the works above remain largely hand-crafted, other methods [omrvcen2009autonomous, clavera2017policy] explore the model-free planning of pushing motions to move objects to target positions that are more favorable for pre-designed grasping algorithms. Boularias et al. [boularias2015learning] explores the use of reinforcement learning for training control policies to select among push and grasp proposals represented by hand-crafted features. However, their method models perception and control policies separately (not end-to-end); it relies on model-based simulation to predict the motion of pushed objects and to infer its benefits for future grasping. Similarly, Yu et al. [yu2023novel] also separates push planning and grasp planning by training two networks. More recently, some works [zeng2018learning, yang2021collaborative, deng2019deep, tang2021learning] have been training model-free end-to-end policy using reinforcement learning to push and separate objects in clutter for better grasping. Our method is also model-free and trained end-to-end, making no assumptions about the shapes or dynamics of the objects. However, unlike all the previous works cited here, ours is the first to demonstrate a pushing behavior for grasping inside granular media.
@@
Active Tactile Exploration. Tactile sensors need an intelligent exploration policy to actively interact with the target object and efficiently gather sparse and local tactile readings. These exploration policies can be roughly divided into heuristics-based and learning-based. One of the most popular heuristics-based exploration policies is contour-following [martinez2013active, yu2015shape, suresh2021tactile]. Other heuristics include information gain (uncertainty reduction) [hebert2013next, xu2013tactile, schneider2009object, driess2017active], attention cubes [rajeswar2021touch], Monte Carlo tree search [zhang2017active], and dynamic potential fields [bierbaum2009grasp]. Some works study tactile exploration policy specifically for decision making, such as object identification, where they pre-train a discriminator with a pre-collected dataset and then use it to estimate action quality with Bayesian methods to reduce uncertainty [fishel2012bayesian, lepora2013active, martinez2017active, kaboli2017tactile, kaboli2019tactile]. Depending on the task, heuristic-based exploration policies can improve efficiency and require no training, but their performance can deteriorate significantly when sensor noise shows up.
@@
In contrast, Xu et al. [xu2022tandem, xu2023tandem3d] study learning-based exploration policy trained through reinforcement learning in an unsupervised fashion. They also aim for object identification, and their method has two separate components, trained jointly: a discriminator for decision-making and an explorer for guiding the exploration. Different from these previous works, our method aims to grasp buried objects inside granular media, which is a more challenging and novel task. In addition, our method only has one end-to-end policy that guides both the pushing and grasping actions.
3 Method
@@
We formulate our problem as follows. A robotic manipulator is presented with a workspace filled with granular media. An object is buried completely under the granular media, placed with unknown orientation or exact position, such that the object is in the way of the gripper and a contact is guaranteed if the gripper moves straight forward. The goal of the manipulator is to retrieve the buried object, by lifting it clear of the granular environment. We assume that the object can start out fully buried with no part visible from the surface. Our goal is to perform the task without relying on vision and using only touch sensing.
@@
Tactile sensing in granular media requires the ability to distinguish between readings due to ubiquitous contacts with the granules themselves, and readings indicating contact with the target object. As explained later in this section, we use a simple filtering technique to make this distinction. However, the nature of operation in such environments increases the noise inherent in any tactile sensor. We thus need an exploration policy that is well-suited for handling tactile noise and spurious or erroneous readings.
@@
3.1 Action Space Definition and Pushing Behavior
We formulate our problem as a model-free reinforcement learning task. Object exploration and retrieval are achieved through a series of discrete translation/rotation actions of the gripper, where the translation action is two-dimensional on a 2D $xy$ plane, and the rotation action is one-dimensional along the $z$ axis. At each step, our policy takes in a sequence of contact locations and forces from the past steps and then outputs the next action for the gripper to take for exploration. Our policy also determines when to stop pushing and execute the grasp.
@@ 3.1 Action Space Definition and Pushing Behavior
Specifically, our action primitives consist of a planar pushing action and a grasping action. The gripper moves in a 2D plane and is always perpendicular to the table, as shown in Fig. 1. Given the current gripper pose (position $[x,y]$ and rotation $\theta$), the action is a 4-dimensional vector $dx,dy,d\theta,G]$. A value of $G=0$ indicates a pushing action, in which case $dx,dy,d\theta$ indicate the change to $x$, $y$, and $\theta$, respectively. For $x,y$, the change is 1cm and for $\theta$, the change is $15^{\circ}$; namely, $dx,dy\in\{-1\text{cm},0,+1\text{cm}\}$ and $d\theta\in\{-15^{\circ},0,+15^{\circ}\}$. Notice that in this action space formulation, it is possible to change $x,y$, and $\theta$ simultaneously in one action, by setting the corresponding elements at the same time. $G=1$ corresponds to a grasping action. In this case, the values of $dx,dy,d\theta$ are ignored, the gripper is closed, and the object is lifted. The episode is terminated when a grasping action is executed or a predefined maximum number of steps is reached.
As discussed in the results section, this formulation results in the learned policy effectively making use of sequences of pushing behaviors of various lengths. However, a critical aspect of learning such behaviors is training in the presence of noise, which we discuss later in Sec. 4.3.
3.2 Observation Space and Tactile Data Processing
@@
Figure 3: Curriculum training strategy. The numbering of the frames corresponds to their order within a single episode. We first pre-train our policy on a tabletop environment and then fine-tune it inside granular media. We then zero-shot deploy our policy on the real hardware for evaluation. This curriculum allows the policy to converge to a higher performance within a shorter training time.
At each time step $t$, the observation of our policy is a sequence of contact locations $p_{t},p_{t-1},\dots,p_{t-h}$, net forces $f_{t},f_{t-1},\dots,f_{t-h}$, and previous actions $a_{t-1},a_{t-2},\dots,a_{t-h-1}$ from the last $h=10$ steps. Observations from both fingers are concatenated to form the final observation of the policy. Both the contact location $p_{t}$ and the net force $f_{t}$ are 3-dimensional vectors under the finger’s frame of reference (using the center of the finger base as the origin). The previous action $a_{t-1}$ is a 4-dimensional vector, as discussed in Sec. 3.1. When there is no contact recorded after a particular step, we simply fill in zeros in the location and force vectors. Compared to high-dimensional or multimodal tactile data, contact locations and forces are easy to simulate, enabling a large amount of training in simulation and zero-shot transfer to real-world experiments.
@@
Tactile observation in granular media is much more noisy compared to open air. First, there might not be direct contact between the finger and the target object when pushing contacts are detected. Instead, there could be a layer of granules between the finger and the target object. This layer can potentially introduce more noise to the force directions. Second, for small and light objects, pushing contact can be hard to distinguish from ubiquitous contacts, and, depending on how fast the finger accelerates and moves inside the medium, pushing contacts can be falsely reported even though the finger has not started pushing the target object.
Ideally, object- and granular media-specific force filtering thresholds might be needed for better filtering accuracy. However, in this work, the filtering threshold only differs between simulation and real hardware but remains the same for all objects, and we observe that our policy learns to be robust to spurious pushing contacts. The filtering threshold is found experimentally through calibration. During the calibration phase, we drag the submerged finger through the granular media without any buried objects. The dragging follows a predefined trajectory with multiple changes in direction. We record the maximum force sensed during the whole trajectory. We then find the minimum integer value larger than the maximum force to be the chosen force threshold. We do the calibration process separately in simulation and on the real hardware, giving us different filtering thresholds.
3.3 Reward Design and Policy Architecture